OpenStudy (anonymous):

Solve 2square root of x plus seven. - 10 = -8

OpenStudy (btaylor):

\[2\sqrt{x+7}-10=-8\] Is that the right equation? If so, first, Isolate the radical (add ten to both sides, divide by two)

OpenStudy (anonymous):

And Then What ?

OpenStudy (anonymous):

can you tell me step by step ?

OpenStudy (btaylor):

1. Add ten to both sides \[2\sqrt{x+7}=2\] 2. Divide both sides by two. \[\sqrt{x+7}=1\] 3. Square both sides \[x+7 = 1\] 4. Subtract 7 from both sides \[x=1-7=-6\] Then plug -6 in for x in the original equation, and you get the right answer.

OpenStudy (anonymous):

Thank You So Much